OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 288 cu. in.

OpenStudy (espex):

The volume of a sphere is:\[v=\frac{4}{3} \pi r^{3}\] and the area of a sphere is: \[A=4\pi r^{2}\]So \[288=\frac{4}{3} \times \pi \times r^{3}\] Solve equation 1 for r and plug that value into the second equation to find the surface area.